OpenStudy (anonymous):

Heeeeellllppp Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11

OpenStudy (anonymous):

isolate for a variable in one equation and sub it into the other: a + b = 4 so b = 4 - a (now sub this into second given equation) 2a + 3 (4 - a) = 11 (now simplify and isolate for a) 2a + 12 - 3a = 11 1 = 1a so a = 1 (now plug a back into first equation to get b) a + b = 4 (1) + b = 4 so b = 3

OpenStudy (anonymous):

Awesome thank you soo much.